0. Doug Bernard specializes in cross-rate arbitrage. He notices the following quotes. Swiss franc/dollar = SFr1.5971/$ AustraCross-Rate Arbitrage
Answer #1

The cross-rate between Australian dollars and Swiss franc is A$/SFr = A$/$ x $/SFr = (A$/$)/(SFr/$) = 1.8215/1.5971 = 1.1405.

But, given cross-rate is higher A$1.1.1440/SFr than what we calculate implicitly.

So, arbitrage opportunity exists.

From given cross-rate of A$1.1440/SFr, 1 Swiss franc is worth A$1.1440, whereas the cross-rate based on the calculated rate implies that 1 Swiss franc is worth A$1.1405.

Thus, the Swiss franc is overvalued relative to the A$ in the given cross-rate, and Doug Bernard’s strategy for arbitrage should be to sell Swiss francs and buy A$ as per the given cross-rate.

Sell dollars to get Swiss francs Sell $1,000,000 to get $1,000,000 x SFr1.5971/$ = SFr1,597,100.

Then sell the Swiss francs to buy Australian dollars SFr1,597,100 to buy SFr1,597,100 x A$1.1440/SFr = A$1,827,082.40.

Sell Australian dollars for dollars  A$1,827,082.40 for A$1,827,082.40/A$1.8215/$ = $1,003,064.73.

Hence arbitrage profit is $1,003,064.73 - $1,000,000 = $3,064.73.